(1 point) Solve the system of equations. S x = y - 7 8x - y = 0 help (fractions) y = help (fractions)
Answer #1

Somtion au Solve the system of equations B2-y= 0 2 = t-h e x-y=- solue equation x-y=- and 2 8x-ly- = Ze 17x=772 12=1 put x=1